OpenStudy (anonymous):

Find the next number in the series.. 3,7,32,211, ?, ?

OpenStudy (anonymous):

another question???‎ 976 : 684 375 : 225 864 : 512 597 : ?

OpenStudy (anonymous):

wats the ansr of previous quesn............???

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

sorry yaar

OpenStudy (anonymous):

dont kno

OpenStudy (dumbcow):

theres no relationship that i can see

OpenStudy (anonymous):

there is relationship .

OpenStudy (dumbcow):

next 2 numbers are 3796 and 45000

OpenStudy (anonymous):

would u plz explain?

OpenStudy (dumbcow):

its an educated guess like i said i dont see a constant relationship

OpenStudy (dumbcow):

using logs i found 3^1.77 = 7 7^1.78 = 32 32^1.54 = 211 so i used 211^1.54 = 3796 3796^1.3 = 45000

OpenStudy (anonymous):

im here, look at the name of the post

OpenStudy (anonymous):

notice you write 32, not 31